OpenStudy (anonymous):

Suppose you pick 4 cards randomly from a well-shuffled standard deck of 52 playing cards. The probability that you draw the 2, 4, 6, and 8 of spades in that order is

OpenStudy (kropot72):

If the cards are drawn without replacement, the probability is given by: \[\large \frac{1}{52}\times\frac{1}{51}\times\frac{1}{50}\times\frac{1}{49}\]

OpenStudy (anonymous):

thank you! :)

OpenStudy (kropot72):

You're welcome :)